OpenStudy (anonymous):

[8.04] What are the factors of 4x2 + 13x + 3? (4x + 3)(x + 1) (4x + 1)(x + 3) (2x + 3)(2x + 1) (2x − 1)(2x − 3)

OpenStudy (anonymous):

solution of equation are -1/4 and -3 4x^2+13x+3=4(x+1/4)(x+3)=(4x+1)(x+3)